Three consevutive odd integers are such that the sum of the first and second is 31 less than 3 times the third. Find the integers.
Answer by edjones(8007)

Let the integers be x, x+2, x+4
x+x+2+31=3(x+4)
2x+33=3x+12
x=21
23 and 25 are the others.
.
Ed